An Inessential^{1} Quick Reference to Athena

1 Introduction

This short document was prepared by the Student Information Processing Board (SIPB) at MIT. The SIPB is a volunteer group of students who help computer users at MIT and work on computer-related projects. Our office is on the fifth floor of the Student Center (W20-557), next to the Athena cluster, and our phone number is x3-7788. Please stop by or call if you have any questions.

This is not even close to a complete guide to Athena. This document is simply the least you need to know to use Athena. The best source of further information is online at http://web.mit.edu/olc/ and http://stuff.mit.edu/sipb/docs/. Printed copies of these documents are also avaialble from CopyTech and the SIPB office.

2 Logging In and Out

When you type your username and password in the login window, you will be “logged in” to Athena. While you are logged in you can read and send e-mail, write papers, play games, browse the web, and lots more. When you are done with your session, left-click the mouse on the “EXIT” button at the bottom of the screen to log out. Don’t turn the machine off; once you have logged out, the workstation is ready for the next user, and the computer no longer has access to your data and personal records.

The large window on the bottom-right side of your screen is a “terminal.” This is a program that allows you to enter commands under Athena. You can start another terminal by clicking the athena% icon near the bottom-left of the screen. You can have as many terminals as you want.

3 Your Password

If you want to change your password, click inside the terminal and type the line passwd at the athena% prompt.

You can never be too careful with your password. Since you are legally responsible for anything done with your account, and your password lets you access your confidential records, use a password that is hard for humans and computers to guess, do not use the same password elsewhere, and log out when you are done with a computer.

4 Directories and Lockers

Your information is stored in files, which are stored in directories. You can also have subdirectories within other directories to keep your files organized.

The locker system allows you to use the same directories on all Athena workstations. Lockers exist for all users as well as most organizations. When you attach a locker, it will be available inside the directory /mit. Your own locker, also known as your “home directory”, is named after your username (e.g., /mit/jruser). Because of the locker system, your home directory and therefore your files can be accessed on any workstation, and none of your personal files are stored on the workstation itself.

To use programs in a locker, type add followed by the locker name at the athena% prompt. For example, you can type add matlab and then matlab (on the next line) to run Matlab. If you just want to use files, you can use attach instead. For example, typing attach shakespeare will let you access the Bard’s plays

and sonnets under /mit/shakespeare.

To change your current directory, type cd followed by the name of the directory you want to go to (for example, cd /mit/shakespeare. If you just type cd by itself, you’ll go back to your home directory.

5 Subdirectories

You can create subdirectories by typing mkdir followed by the name you want to call the directory (e.g., mkdir funstuff). You can move into this directory with the cd (Change Directory) command.

By default, people can list filenames but not view the files themselves. In the Private directory, others cannot list files; in the Public and www^{2} directories, everyone can view the files. Another useful directory is OldFiles, which contains a backup of yesterday’s files. All of these subdirectories are in your locker automatically.

1SIPB Documents are traditionally called “Inessential,” even when they contain useful information.

2Any files in your www directory are also available on the web at http://www.mit.edu/ username.

6 Listing, Copying, and Deleting (Removing) Files

The easiest way to view files is using “Nautilus”, the GNOME file browser. Double-click the icon labeled “your username’s Home” at the top-left of your screen. This will bring up a window with icons for all the files in your home directory. Double-click on an icon to open the file or directory. You can move files by dragging them (hold down the mouse button). To delete a file, drag it to the Trash icon on the desktop. You can undelete a file by dragging it back.

If you want to use the Athena prompt, which is sometimes faster, you can list the files with ls (or ls -l for more detail). To copy a file from one name to another, type cp followed by the old name and the new name (e.g., cp paper paper-backup). When you want to get rid of a file, you can use the delete command (e.g., delete junk). If you made a mistake, the undelete command will bring your file back if you do it within around 3 days (e.g., undelete junk).

You have a quota of one gigabyte (about a million kilobytes) of data. If you receive a message saying you are nearing your quota, you will not be able to save any new large files, and may have trouble logging in. If you are nearing your quota, you will need to use rm to delete files permanently. You cannot undelete a file deleted with rm. You should also right-click on the Trash icon and click “Empty Trash.”

7 Web Browsing

You can use the Firefox browser on Athena by clicking the “WWW” icon at the bottom left. Click the button only once; it may take up to a minute to start on some computers.

For most of the websites at MIT, instead of logging in, you use an “MIT certificate”. You can get one by going to web.mit.edu/ist/cert and clicking on both “Get” links. This will securely store your certificate inside your Athena account. (Note that Firefox will ask you to create a master password, to encrypt your certificate, after you tell MIT your Kerberos password. You should avoid using the same password.)

8 Mail

If you have mail, you will see a message in the “Console” window when you log in. The recommended way to check your e-mail is with Evolution, which you can start with the Mail button at the bottom-left. Evolution works similarly to mail clients you may have used before.

Two other ways are Webmail and Pine. Although Webmail is not as powerful as the other two, you can use it from any computer (as well as Athena) by visiting http://webmail.mit.edu/. If you prefer the command line, you may like Pine, which you can use by typing pine at a terminal.

9 Creating and editing documents

To view or edit plain-text files on Athena, the easiest option is to type gedit, followed by the name of the file, followed by an ampersand (e.g., gedit filename &). You need the ampersand on most Athena programs that bring up a window, or you will not get your athena% prompt back until you close the program.

Another tool for editing plain text files is Emacs, which has a bit of a learning curve but is far more powerful than gedit. To get started with Emacs, click the Emacs button at the bottom of the screen, and then open the tutorial from the Help menu.

For formatted documents, including word processing, spreadsheets, and presentations, you can use the OpenOffice.org package. This software is compatible with Microsoft Office files as well as the OpenDocument format. You can run OpenOffice.org by clicking the footprint icon labeled “Menu”, at the bottom left, and selecting “Editing/Publishing”, “Office Productivity Suites”, and “Open Office”.

When you are writing theses or scientific papers, most departments will want you to use L^{A}TEXinstead. For more information on using L^{A}TEX, read SIPB’s “Inessential Guide to L^{A}TEX”, available online or from our office.

10 Learning More about Athena

The following resources are available:

Resource  How to Use
Pocket Guide to Athena  Athena Documentation

Type help at the athena% prompt or visit http://web.mit.edu/olc/.

On-Line Consultants  Type olc at the athena% prompt.
Athena Consultants  Call x3-4435 or stop by N42.
Athena Accounts  Call x3-1325 or visit N42 for help with accounts or passwords.
Manual Pages  Type man and the command you want to look up (e.g., man ls).
SIPB  Call x3-7788, stop by W20-557, or e-mail sipb@mit.edu.

You can also ask other users for help online, by using Zephyr. Type zctl sub help to join the “help” class, and then use zwrite -c help -i topic (where topic is a one-word description, such as emacs) to ask your questions. Although this is an unofficial service with no guarantees, chances are you will get a correct answer.

Version 1.3 by amgreene 18-Oct-89; Version 1.4 by amgreene 21-Aug-90; Version 1.5 by sorokin 3-Oct-93; Version 2.0 by geofft 7-Oct-06.